Can you please help me factor 2t^2 - 19 - 6t =
Answer by checkley71(8403)

2t^2-19-6t=0
2t^2-6t-19=0 using the quadratic equation we get
x=(-b+-sqrt[b^2-4ac])/2a
x=(6+-sqrt[36-4*2*-19])/2*2
x=(6+-sqrt[36+152])/4
x=(6+-sqrt188])/4
x=(6+-13.71)/4
x=(6+13.71)/4
x=19.71/4
x=4.93 solution
x=(6-13.71)/4
x=-7.71/4
x=-1.93 solution